OpenStudy (anonymous):

. A rope pulling a sled is sloped at 25° upward at the point of contact. The tension in the rope is 40.0 N. What is the horizontal component of the force? please can anyone help? this is my first time here ..

OpenStudy (anonymous):

okay, so, we split this force into its horizontal components and its vertical components through the use of some trigonometrics. \[cos\theta = \frac{adjacent}{hypotenuse}\] The adjacent side in this case being the horrizontal component, the hypotenuse is the 40.0 N tension in the rope and theta there is that lovely little 25 degree angle that we have. So from here its a simple matter of rearranging the equation to find Adjacent on its own: \[hyp*cos\theta = Adj=36.25 N\]

OpenStudy (anonymous):

thanks so much :)